OpenStudy (anonymous):

simplify the radical expression square root of 250h^4k^5

OpenStudy (anonymous):

is this algebra or what tipe of math

OpenStudy (anonymous):

Yeah

OpenStudy (mertsj):

\[\sqrt{250h^4k^5}=\sqrt{25(10)h^4k^4k}=5h^2k^2\sqrt{10k}\]